# CS-Bench: A Comprehensive Benchmark for Large Language Models towards Computer Science Mastery

### v1

@@ C.1 Detailed Design Motivation and Statistics of CS-Bench
Overview Concepts and elements of data structure, Temporal and spatial complexity… DSA 84
Linear List Linear tables, Sequential tables and Linked lists… DSA 138
Stack, Queue,and Array Shared stack, Circle queue, Arrays,Special matrices… DSA 176
String Concept and operation of strings, Pattern matching of strings… DSA 66
Tree Binary trees, Traversal of trees ans forests, Huffman tree… DSA 214
Graph Concepts of graphs, Traversals of graphs,Application of graphs… DSA 184
Searching Sequential search, Half-split search, Chunked search, Red-black tree, B-tree and B+ tree, Hash search… DSA 158
Sorting Insert Sorting, Swap Sorting, Selection Sorting, Merge Sorting, Heap Sorting, Merge Sorting, Cardinality Sorting, External Sorting Algorithms… DSA 178
Overview Hardware and performance indicators of computers… CO 112
Data Representation and Operation Number system and encoding, Representation and operation of fixed-point nu- mbers and floating-point numbers… CO 218
Storage System Main Memory, External Memory, Cache Memory, Virtual Memory… CO 224
Instruction System Instruction format, Instruction addressing method, CISC and RISC… CO 156
Central Processing Unit Functions of CPU, Instruction execution process, CPU internal bus and data path, CPU hard wiring design and micro programming, Exception and inter- rupt mechanisms, Instruction pipelines, and multiprocessor concepts… CO 244
Bus Overview of the bus, Bus arbitration, Bus operation and timing, Bus standards… CO 134
Input/Output System I/O interfaces and methods… CO 156
Overview and Architecture Concepts, compositions, functions of computer networks, Architecture and reference models of computer networks… CN 148
Physical Layer Fundamentals of Communication Theory, Transmission Media and Physical Layer Devices… CN 164
Data Link Layer Data frames, Error control, Flow control and Reliable transmission, Media acc- ess control, Local and wide area networks, and data link layer devices… CN 316
Network Layer Overview of network layer functions, Routing algorithms, IPv4 and IPv6, Rou- ting protocols, IP multicast, Mobile IP, Router… CN 300
Transport Layer The services provided by the transport layer, UDP and TCP protocols… CN 182
Application Layer Network application model, Domain name system DNS, FTP protocol, World Wide Web, and HTTP… CN 204
Overview Concepts of operating systems, Development and classification of operating systems, Operational mechanisms and architecture of operating systems… OS 166
Processes and Threads Processes and threads, Scheduling of processors, Synchronization and mutual exclusion of processes, Deadlock issues… OS 350
Memory Management Concept of memory management, Concept of virtual memory management, and methods of virtual memory management… OS 216
File Management File systems, Organization and management of disks… OS 166
Input/Output Management I/O devices and control methods, I/O core subsystem, Buffer management… OS 184
@@
To assess the ability of LLMs in addressing CS problems in various linguistic environments, and to adapt CS-Bench for the evaluation of a wider range of LLMs, CS-Bench comprises bilingual Chinese-English data, with each language accounting for 50%. The English data is obtained through translation by GPT-4, followed by manual verification of processed Chinese data.
C.2 Distribution of Word Lengths
Due to CS-Bench containing both English and Chinese languages, we separately compute the distributions of word lengths for questions and answers in CS-Bench (English) and CS-Bench (Chinese) across various task formats, as illustrated in Figure 8 and Figure 9. For Multiple-Choice questions, the question length includes both the question itself and the four options. Since Multiple-Choice and Assertion questions are comprehension tasks, the answers consist of only one character (A/B/C/D or True/False). For generation tasks, Fill-in-the-blank answers are relatively short, with an average word length of approximately 2, whereas Open-ended questions typically yield longer answers as they entail detailed explanatory processes.
@@ C.2 Distribution of Word Lengths
Figure 9: Question and answer lengths of each task format in CS-Bench (Chinese).
C.3 CS-Bench Examples
We present samples from various domains in Table 6, samples of different task formats in Table 7, samples of knowledge and reasoning types in Table 8, and samples from different languages in Table 9.
@@ C.3 CS-Bench Examples
Table 8: Examples of knowledge-type and reasoning-type.
Type Example
@@
English Question: For a linear list with sequential storage, the operation with a time complexity of O(1) should be (). A: Sort n elements in ascending order. B: Remove the i-th ($1\leq i\leq n$) element. C: Change the value of the i-th element ($1\leq i\leq n$). D: Insert a new element after the i-th ($1\leq i\leq n$) element. Answer: C Analysis: The time complexity for sorting n elements is at least O(n) (when initially ordered), and typically $\mathcal{O}(n\log_{2}n)$ or $\mathcal{O}(n^{2})$. Options B and D are clearly incorrect. Sequential lists support random access by index.
@@ Appendix D More Details on Experiment Setup
In D.1, we present the question templates used to prompt models for each type of task. In D.2, we show the prompts used for GPT-4 to score models’ answers to fill-in-the-blank and open-ended questions, and validate the effectiveness of GPT-4’s automatic scoring through consistency experiments with human scoring. In D.3, we detail the experimental environment used to implement model inference. In D.4, we introduce all the evaluated model families.
@@ D.2 Details of GPT-4 Scoring
In Table 12, we present the prompts utilized to instruct GPT-4 in scoring the outputs of LLMs in CS generation tasks, encompassing both Fill-in-the-blank and Open-ended questions.
@@
To assess the effectiveness of GPT-4 scoring in evaluating LLM responses, we conduct a consistency experiment between GPT-4 prediction scores and manual scores. For Fill-in-the-blank and Open-ended types, we randomly sample 100 instances from the GPT-4 scoring samples and employ three human annotators to score these predicted results. In Table 11, we report the consistency scores among human annotators (measured by Cronbach’s alpha), as well as the consistency scores between the average human annotation scores and GPT-4 scoring (measured by Pearson correlation coefficient). The excellent consistency between human and GPT-4 scores validates the effectiveness of GPT-4 scoring.
@@ D.3 Details of Inference Implementation
For all open-source models, we utilize a cluster with 8 NVIDIA A100-80GB GPUs to run the inference, and we use vLLM [78] for inference acceleration, applying the corresponding chat templates and the same hyper-parameters: batch size=1, temperature=0, top-p=1.0, and max_tokens=2048. For all closed-source models with API access, we also adopt the generation scheme with temperature=0, and simply run the inference with CPUs, which typically completes within a day. During the evaluation of GPT-4, we also applied the setting of temperature=0. To avoid error bias, we conducted the experiments 3 times and took the average of the scores. For models supporting web search or tool calls, we disable these features to ensure a fair comparison.
@@ D.4 Details of the Models being Evaluated
[37] is a family of lightweight, open models from Google, built from the same research and technology used to create the Gemini models. They are text-to-text, decoder-only large language models, available in English, with open weights, pre-trained variants, and instruction-tuned variants. The Gemma model excels on academic benchmarks in language understanding, reasoning, and security. Gemma publishes models in two sizes (2 billion and 7 billion parameters) .
@@
[38] is an upgraded version of Llama developed by MetaAI. It utilizes more robust data cleaning and mixing techniques, and up-samples sources closest to factual information, which can enhance knowledge and reduce hallucinations. Additionally, it employs Grouped-Query Attention technology to lessen reliance on memory.
@@
[39] is the latest generation of large language models developed by MetaAI. The training dataset for Llama 3 is seven times larger than that used for Llama 2, with the amount of code included being four times that of Llama 2. Compared to previous versions of the model, it has seen a tremendous enhancement in reasoning, code generation, and instruction following capabilities.
@@
[79] is an instruction-tuned language model for Chinese and English users with various abilities such as roleplaying and tool-using built upon the Meta-Llama-3-8B-Instruct model.
@@
[80] is a next-generation conversational pre-trained model jointly released by Zhipu AI and KEG Lab of Tsinghua University. ChatGLM3-6B adopts a newly designed Prompt format, in addition to regular multi-turn dialogue. It also natively supports complex scenarios such as function call, code interpretation.
@@
[41] is a large-scale multilingual model developed by Baichuan Company. It adopts several advanced techniques in its design and training process, including Rotary Position Embedding, a novel position encoding technique, SwiGLU activation function, and memory efficient attention mechanism. Compared with Baichuan1, its performance has been greatly improved.
@@
[42] is an open-source large-scale language model developed by Shanghai AI Laboratory. This model has good processing ability for ultra long texts and adopts COOL RLHF technology. It solves human preference conflicts through a conditional reward model and performs multiple rounds of online RLHF to improve the model’s alignment ability.
@@
[81] is a family of language models developed by Alibaba. It has features such as SwiGLU activation, attention QKV bias, group query attention, mixture of sliding window attention and full attention, etc. Qwen 1.5 series models have strong basic capabilities including language understanding.
@@
[44], a 7-billion-parameter language model designed for superior performance and efficiency, which is developed by Mistral AI. Mistral 7B leverages Packet Query Attention (GQA) for faster inference, combined with Sliding Window Attention (SWA) to efficiently process sequences of arbitrary length while reducing inference costs.
@@
[45] is a Sparse Mixture of Experts (SMoE) language model developed by Mistral AI. Its architecture is the same as that of the Mistral 7B, except that each layer consists of 8 feedforward blocks (i.e., experts). Mixtral has demonstrated exceptional abilities in math, code generation, and tasks that require multilingual understanding.
@@
[46] is a family of models released by DeepSeek-AI, and its core architecture borrows from the Llama model. This family of models employs Multi-Head Attention (MHA) and Group Query Attention (GQA) techniques, which significantly enhance their performance and efficiency. Furthermore, DeepSeekLLM demonstrates strong bilingual capabilities in both Chinese and English.
@@
[47] is the higher-performance successor to PaLM released by Google, which differs in terms of dataset mixing. Compared to the first-generation PaLM version, it uses a smaller model but performs more training calculations. It also relies on more diverse pre-training targets.
Claude
Claude2.1[48] and Claude3 [49] are AI models developed by Anthropic, showcasing advanced language understanding and generation capabilities. Utilizing the constitutional AI framework, Claude models are designed to ensure helpfulness and trustworthiness.
@@
GPT-3.5 [5], GPT-4 [50] and GPT-4o [6], released by OpenAI, are part of the GPT-series models enhanced by a three-stage reinforcement learning with human feedback (RLHF) algorithm. This algorithm not only improves the models’ ability to follow instructions but also significantly reduces the generation of harmful or toxic content. Additionally, GPT-4 supports image inputs and achieves human-level performance on various benchmarks. GPT-4o, the latest model developed by OpenAI, boasts powerful real-time reasoning, language interaction, and multimodal capabilities.
@@
[82] is a new generation base large model developed by Zhipu AI. It has strong tool calling and multi-modal capabilities, as well as strong mathematical reasoning ability and code generation ability.
@@
[83] ERNIE3.5 and ERNIE4 are large language models developed by Baidu. ERNIE3.5 is capable of processing text data in multiple languages and has a good understanding and representation ability for entities and relationships in text. Ernie 4 has adopted more advanced knowledge graph information and more advanced knowledge integration technology, further improving the performance of the model.

### v2

@@ C.2 Detailed Design Motivation and Statistics of CS-Bench
Overview Concepts and elements of data structure, Temporal and spatial complexity… DSA 168
Linear List Linear tables, Sequential tables and Linked lists… DSA 276
Stack, Queue,and Array Shared stack, Circle queue, Arrays,Special matrices… DSA 352
String Concept and operation of strings, Pattern matching of strings… DSA 132
Tree Binary trees, Traversal of trees ans forests, Huffman tree… DSA 428
Graph Concepts of graphs, Traversals of graphs,Application of graphs… DSA 368
Searching Sequential search, Half-split search, Chunked search, Red-black tree, B-tree and B+ tree, Hash search… DSA 316
Sorting Insert Sorting, Swap Sorting, Selection Sorting, Merge Sorting, Heap Sorting, Merge Sorting, Cardinality Sorting, External Sorting Algorithms… DSA 356
Overview Hardware and performance indicators of computers… CO 224
Data Representation and Operation Number system and encoding, Representation and operation of fixed-point nu- mbers and floating-point numbers… CO 436
Storage System Main Memory, External Memory, Cache Memory, Virtual Memory… CO 448
Instruction System Instruction format, Instruction addressing method, CISC and RISC… CO 312
Central Processing Unit Functions of CPU, Instruction execution process, CPU internal bus and data path, CPU hard wiring design and micro programming, Exception and inter- rupt mechanisms, Instruction pipelines, and multiprocessor concepts… CO 488
Bus Overview of the bus, Bus arbitration, Bus operation and timing, Bus standards… CO 268
Input/Output System I/O interfaces and methods… CO 312
Overview and Architecture Concepts, compositions, functions of computer networks, Architecture and reference models of computer networks… CN 296
Physical Layer Fundamentals of Communication Theory, Transmission Media and Physical Layer Devices… CN 328
Data Link Layer Data frames, Error control, Flow control and Reliable transmission, Media acc- ess control, Local and wide area networks, and data link layer devices… CN 632
Network Layer Overview of network layer functions, Routing algorithms, IPv4 and IPv6, Rou- ting protocols, IP multicast, Mobile IP, Router… CN 600
Transport Layer The services provided by the transport layer, UDP and TCP protocols… CN 364
Application Layer Network application model, Domain name system DNS, FTP protocol, World Wide Web, and HTTP… CN 408
Overview Concepts of operating systems, Development and classification of operating systems, Operational mechanisms and architecture of operating systems… OS 332
Processes and Threads Processes and threads, Scheduling of processors, Synchronization and mutual exclusion of processes, Deadlock issues… OS 700
Memory Management Concept of memory management, Concept of virtual memory management, and methods of virtual memory management… OS 432
File Management File systems, Organization and management of disks… OS 332
Input/Output Management I/O devices and control methods, I/O core subsystem, Buffer management… OS 368
@@
To assess the ability of LLMs in addressing CS problems in various linguistic environments, and to adapt CS-Bench for the evaluation of a wider range of LLMs, CS-Bench comprises quadrilingual (English, Chinese, French, and German) data, with each language accounting for 25%. The other data is obtained through translation by GPT-4, followed by manual verification of processed Chinese data.
C.3 Distribution of Word Lengths
We compute the distributions of word lengths for questions and answers in CS-Bench (English) across various task formats, as illustrated in Figure 11. For Multiple-Choice questions, the question length includes both the question itself and the four options. Since Multiple-Choice and Assertion questions are comprehension tasks, the answers consist of only one character (A/B/C/D or True/False). For generation tasks, Fill-in-the-blank answers are relatively short, with an average word length of approximately 2, whereas Open-ended questions typically yield longer answers as they entail detailed explanatory processes.
@@ C.3 Distribution of Word Lengths
C.4 CS-Bench Examples
We present samples of knowledge and reasoning types in Table 6, samples from various domains in Table 7, samples of different task formats in Table 8, and samples from different languages in Table C.4.
Table 6: Examples of knowledge-type and reasoning-type.
Type Example
@@ C.4 CS-Bench Examples
English Question: For a linear list with sequential storage, the operation with a time complexity of $O(1)$ should be (). A: Sort n elements in ascending order. B: Remove the i-th ($1\leq i\leq n$) element. C: Change the value of the i-th element ($1\leq i\leq n$). D: Insert a new element after the i-th ($1\leq i\leq n$) element. Answer: C Analysis: The time complexity for sorting n elements is at least O(n) (when initially ordered), and typically $O(n\log_{2}n)$ or $O(n^{2})$. Options B and D are clearly incorrect. Sequential lists support random access by index.
@@
Question: Pour une liste linéaire stockée de manière séquentielle, l’opération dont la complexité temporelle algorithmique est $O(1)$ devrait être (). A: {CJK*}UTF8gbsnTrier n éléments dans l’ordre croissant. B: Supprimer le $i$-ème ($1\leq i\leq n$) élément. C: {CJK*}UTF8gbsnModifier la valeur du ième élément ($1\leq i\leq n$). D: {CJK*}UTF8gbsnInsérez un nouvel élément après le i-ème ($1\leq i\leq n$) élément. Answer: C Analysis: {CJK*}UTF8gbsnLa complexité temporelle pour trier $n$léments est au minimum $O(n)$ (lorsque la liste est initialement ordonnée), généralement $O(n\log_{2}n)$ ou $O(n^{2})$. Les options B et D sont manifestement incorrectes. Une liste séquentielle supporte l’accès aléatoire par indice.
German Question: {CJK*}UTF8gbsnFür sequentiell gespeicherte lineare Listen, welche Operation hat eine algorithmische Zeitkomplexität von $O(1)$? (). A: {CJK*}UTF8gbsnSortiere n Elemente der Größe nach von klein nach groß. B: {CJK*}UTF8gbsnLösche das $i$-te ($1\leq i\leq n$) Element . C: {CJK*}UTF8gbsnÄndere den Wert des i-ten Elements ($1\leq i\leq n$). D: {CJK*}UTF8gbsnFüge nach dem i-ten ($1\leq i\leq n$) Element ein neues Element ein. Answer: C Analysis: Die Zeitkomplexität für das Sortieren von $n$ Elementen beträgt mindestens $O(n)$ (wenn sie anfänglich sortiert sind) und ist normalerweise $O(n\log_{2}n)$ oder $O(n^{2})$. B und D sind offensichtlich falsch. Eine sequentielle Liste unterstützt den zufälligen Zugriff nach Index.
@@ Appendix D More Details on Experiment Setup
In D.1, we present the question templates used to prompt models for each type of task. In D.2, we show the prompts used for GPT-4 to score models’ answers to fill-in-the-blank and open-ended questions, and validate the effectiveness of GPT-4’s automatic scoring through consistency experiments with human scoring. We also explore the effect of different scoring models. In D.3, we detail the experimental environment used to implement model inference. In D.4, we introduce all the evaluated model families.
@@ D.2 Details of GPT-4 Scoring
In Table 13, we present the prompts utilized to instruct GPT-4 in scoring the outputs of LLMs in CS generation tasks, encompassing both FITB and Open-ended questions.
@@
To assess the effectiveness of GPT-4 scoring in evaluating LLM responses, we conduct a consistency experiment between GPT-4 prediction scores and manual scores. For Fill-in-the-blank and Open-ended types, we randomly sample 100 instances from the GPT-4 scoring samples and employ three human annotators to score these predicted results. These three annotators all hold bachelor’s degrees in computer science. Their scoring criteria were consistent with the evaluation standards provided to GPT-4 and can be found in Table 13. In Table 11, we report the consistency scores among human annotators (measured by Cronbach’s alpha), as well as the consistency scores between the average human annotation scores and GPT-4 scoring (measured by Pearson correlation coefficient). The excellent consistency between human and GPT-4 scores validates the effectiveness of GPT-4 scoring.
@@
Effect of Different Scoring Models
To investigate the impact of different models acting as judges, as well as the relationship between the tested model and the judge’s capabilities on the robustness of the evaluation process, we first use the Claude family to construct a set of models with varying abilities: Claude3-Haiku $<$ Claude3-Sonnet $<$ Claude3-Opus $<$ Claude3.5-Sonnet (following the official Claude capability ranking (Anthropic, 2024b)). Next, we pair each two models respectively as the judge and the tested model. The results are as shown in Table 12.
Table 12: Performance of tested models under different scoring models.
Model Claude3-Jaiku (Judge) Claude3-Sonnet (Judge) Claude3-Opus (Judge) Claude3.5-Sonnet (Judge) GPT-4 (Judge)
Claude3-Haiku (Test) 64.61 64.56 63.22 62.45 60.69
Claude3-Sonnet (Test) 64.33 63.75 62.13 62.05 60.17
Claude3-Opus (Test) 71.74 71.29 70.46 69.97 68.63
Claude3.5-Sonnet (Test) 73.97 73.46 72.43 72.27 71.01
Firstly, when all Claude models act as judges, their scoring of the tested models aligns consistently with GPT-4. This effectively demonstrates the robustness of the evaluation process, as consistent ranking results are obtained regardless of whether the tested model’s performance is inferior or superior to that of the judge model (Claude3-Haiku and Claude3.5-Opus representing the two extremes of judge model capabilities). We believe that as long as a model possesses a certain level of discernment, it can serve as a judge, without needing to outperform all tested models. Even the least capable model, Claude3-Haiku, can yield score rankings consistent with GPT-4 during evaluation, while costing only \$0.25 / 1.25 per million tokens for input/output, respectively, significantly reducing evaluation costs. Secondly, we observe that stronger models tend to be more stringent in their scoring. This is reflected in the assessment scores for all tested models, consistently following the order: Claude3-Haiku $<$ Claude3-Sonnet $<$ Claude3-Opus $<$ Claude3.5-Sonnet. We believe this occurs because more capable evaluation models can identify deeper levels of error. Lastly, we find that the models score fairly and do not tend to give themselves disproportionately high scores (as seen when the same model acts as both judge and test model).
@@ D.3 Details of Inference Implementation
For all open-source models, we utilize a cluster with 8 NVIDIA A100-80GB GPUs to run the inference, and we use vLLM (Kwon et al., 2023) for inference acceleration, applying the corresponding chat templates and the same hyper-parameters: batch size=1, temperature=0, top-p=1.0, and max_tokens=2048. For all closed-source models with API access, we also adopt the generation scheme with temperature=0, and simply run the inference with CPUs, which typically completes within a day. During the evaluation of GPT-4, we also applied the setting of temperature=0. To avoid error bias, we conducted the experiments 3 times and took the average of the scores. For models supporting web search or tool calls, we disable these features to ensure a fair comparison.
@@ D.4 Details of the Models being Evaluated
(Team et al., 2024) is a family of lightweight, open models from Google, built from the same research and technology used to create the Gemini models. They are text-to-text, decoder-only large language models, available in English, with open weights, pre-trained variants, and instruction-tuned variants. The Gemma model excels on academic benchmarks in language understanding, reasoning, and security. Gemma publishes models in two sizes (2 billion and 7 billion parameters) .
@@
(Touvron et al., 2023) is an upgraded version of Llama developed by MetaAI. It utilizes more robust data cleaning and mixing techniques, and up-samples sources closest to factual information, which can enhance knowledge and reduce hallucinations. Additionally, it employs Grouped-Query Attention technology to lessen reliance on memory.
@@
(Meta, 2024) is the latest generation of large language models developed by MetaAI. The training dataset for Llama 3 is seven times larger than that used for Llama 2, with the amount of code included being four times that of Llama 2. Compared to previous versions of the model, it has seen a tremendous enhancement in reasoning, code generation, and instruction following capabilities.
Llama3.1-405B
(Dubey et al., 2024) is an auto-regressive language model developed by MetaAI, utilizing an optimized transformer architecture. Trained on over 15 trillion tokens, it demonstrates excellent flexibility, control, and advanced capabilities. Llama3.1-405B is the largest openly available foundation model released by MetaAI.
@@
(Wang & Zheng, 2024) is an instruction-tuned language model for Chinese and English users with various abilities such as roleplaying and tool-using built upon the Meta-Llama-3-8B-Instruct model.
@@
(Zeng et al., 2022) is a next-generation conversational pre-trained model jointly released by Zhipu AI and KEG Lab of Tsinghua University. ChatGLM3-6B adopts a newly designed Prompt format, in addition to regular multi-turn dialogue. It also natively supports complex scenarios such as function call, code interpretation.
@@
(Yang et al., 2023) is a large-scale multilingual model developed by Baichuan Company. It adopts several advanced techniques in its design and training process, including Rotary Position Embedding, a novel position encoding technique, SwiGLU activation function, and memory efficient attention mechanism. Compared with Baichuan1, its performance has been greatly improved.
@@
(Cai et al., 2024) is an open-source large-scale language model developed by Shanghai AI Laboratory. This model has good processing ability for ultra long texts and adopts COOL RLHF technology. It solves human preference conflicts through a conditional reward model and performs multiple rounds of online RLHF to improve the model’s alignment ability.
@@
(Bai et al., 2023) is a family of language models developed by Alibaba. It has features such as SwiGLU activation, attention QKV bias, group query attention, mixture of sliding window attention and full attention, etc. Qwen 1.5 series models have strong basic capabilities including language understanding.
@@
(Jiang et al., 2023), a 7-billion-parameter language model designed for superior performance and efficiency, which is developed by Mistral AI. Mistral 7B leverages Packet Query Attention (GQA) for faster inference, combined with Sliding Window Attention (SWA) to efficiently process sequences of arbitrary length while reducing inference costs.
@@
(Jiang et al., 2024) is a Sparse Mixture of Experts (SMoE) language model developed by Mistral AI. Its architecture is the same as that of the Mistral 7B, except that each layer consists of 8 feedforward blocks (i.e., experts). Mixtral has demonstrated exceptional abilities in math, code generation, and tasks that require multilingual understanding.
Mistral-Large-123B
(MistralAI, 2024) is a language model developed by Mistral AI with a 128K context window, supporting dozens of languages including French, German, Spanish, Italian, Russian, Chinese, Japanese, and Korean, as well as over 80 programming languages. It is designed for single-node inference with long-context applications in mind – its size of 123 billion parameters allows it to run at large throughput on a single node.
@@
(Bi et al., 2024) is a family of models released by DeepSeek-AI, and its core architecture borrows from the Llama model. This family of models employs Multi-Head Attention (MHA) and Group Query Attention (GQA) techniques, which significantly enhance their performance and efficiency. Furthermore, DeepSeekLLM demonstrates strong bilingual capabilities in both Chinese and English.
@@
(Anil et al., 2023) is the higher-performance successor to PaLM released by Google, which differs in terms of dataset mixing. Compared to the first-generation PaLM version, it uses a smaller model but performs more training calculations. It also relies on more diverse pre-training targets.
Claude2
Claude2.1(Anthropic, 2023) and Claude3 (Anthropic, 2024a) are AI models developed by Anthropic, showcasing advanced language understanding and generation capabilities. Utilizing the constitutional AI framework, Claude models are designed to ensure helpfulness and trustworthiness. Claude-3.5-Sonnet (Anthropic, 2024b) is the first product in the Claude-3.5 model series developed by Anthropic. It demonstrates significant improvements in understanding nuance, humor, and complex instructions, and excels at generating high-quality content with a natural, relatable tone. It possesses exceptional coding capabilities, allowing it to independently write, edit, and execute code, along with advanced reasoning and troubleshooting skills. Additionally, it is Anthropic’s most powerful visual model to date.
@@
GPT-3.5 (OpenAI, 2022), GPT-4 (Achiam et al., 2023) and GPT-4o (OpenAI, 2024a), released by OpenAI, are part of the GPT-series models enhanced by a three-stage reinforcement learning with human feedback (RLHF) algorithm. This algorithm not only improves the models’ ability to follow instructions but also significantly reduces the generation of harmful or toxic content. Additionally, GPT-4 supports image inputs and achieves human-level performance on various benchmarks. GPT-4o, the latest model developed by OpenAI, boasts powerful real-time reasoning, language interaction, and multimodal capabilities.
OpenAI-o1
(OpenAI, 2024b) is a new large language model developed by OpenAI, trained with reinforcement learning to handle complex reasoning. It can produce a long internal chain of thought before responding to the user, and o1 refines its chain of thought and improves its strategies by learning to recognize and correct mistakes, break down complex steps into simpler ones, and adopt alternative approaches when the current method fails. It currently has two versions: o1-preview, which has strong reasoning capabilities and broad world knowledge, and o1-mini, a lightweight version with faster reasoning speed.
@@
(AI, 2024) is a new generation base large model developed by Zhipu AI. It has strong tool calling and multi-modal capabilities, as well as strong mathematical reasoning ability and code generation ability.
@@
(Baidu, 2023) ERNIE3.5 and ERNIE4 are large language models developed by Baidu. ERNIE3.5 is capable of processing text data in multiple languages and has a good understanding and representation ability for entities and relationships in text. Ernie 4 has adopted more advanced knowledge graph information and more advanced knowledge integration technology, further improving the performance of the model.